There are important trade-offs involved in preserving, species habitat. One critical

issue is how much habitat to preserve, given the costs of preservation. As an analyst

at the North Carolina Resource Conservation Commission, you have been asked vo

make a recommendation regarding preservation of critical habitat for the purple-

pellied tree trimmer, a bird found only in North Carolina. Fach year, one thousand


Qt =100~P

Where P is the price-per-acre that bird-watchers would be willing to pay to preserve

Qacres of habitat. You are told that the marginal cost of preserving that habitat is

$60,000 per acre.


How many acres of purple-bellied tree trimmer habitat would be preserved in an

economically efficient scenario?

Calculate the net benefits at that efficient quantity.



trimmer, has been discovered in a site


bellied tree A

A new species, the green- . Discovery of this new

i adjacent to the habitat you are proposing to ee ball Fie timer bablat


function for the purp!

pees cane f





1
Expert's answer
2021-09-04T15:18:08-0400

Qt =100 - P, then P = 100 - Q.

The optimal quantity occurs when MR = MC, so:

MR = TR'(Q) = 100 - 2Q,

100 - 2Q = 60,000,

Q = -29,950 units, so at this cost purple-bellied tree trimmer habitat would not be preserved in an economically efficient scenario at all.